OpenStudy (anonymous):

Evaluate the expression for x = two and one fourth. 12x + 6

OpenStudy (anonymous):

For 2 its 30

OpenStudy (anonymous):

for 1/4 its 3

OpenStudy (anonymous):

Hint: \(x=2\frac{1}{4}=2+\frac{1}{4}\)

OpenStudy (anonymous):

Wait, is x two separate quantities, or a fraction?

OpenStudy (anonymous):

it just gave that one info

OpenStudy (anonymous):

i believe 30 is correct though im going to recheck

OpenStudy (smored):

@collegekids 0.25 of 12=3, 3+6=9, not 3

OpenStudy (anonymous):

9 is right

OpenStudy (anonymous):

for 1/4

OpenStudy (anonymous):

thanks @smored

OpenStudy (smored):

no prob